Suppose rRF = 5%, rM = 11%, and rA = 13%.
Calculate Stock A's beta. Round your answer to two decimal places.
If Stock A's beta were 1.2, then what would be A's new required rate of return? Round your answer to two decimal places.
%
Answer #1
Hi,

Please find the calculations as follows:

13 = 5 + b(11 - 5)
13 - 5 = 6b
8 = 6b
b = 1.33

Required rate of return when beta is 1.2

5 + 1.2 (11- 5)
5 + 7.2 = 12.2%

Thanks,
Aman
Answer #2
expected return = risk free rate + beta (market premium)
13 = 5+B(11 -5)
8/6 = B
thus beta = 1.33
if beta = 1.2
required rate of return = 5+1.2(11-5) = 12.2 %